OpenStudy (anonymous):

Connie has saved up $15 to purchase a new CD from the local store. The sales tax in her county is 5% of the sticker price. Write an equation and solve it to determine the value of the highest priced CD Connie can purchase with her $15, including the sales tax. Round your answer to the nearest penny. x + 0.05x = 15; x = $14.29 x - 0.05x = 15; x = $15.79 0.5x = 15; x = $30 5x = 15; x = $3.00

OpenStudy (anonymous):

\[5.00+15.00= 20.00\]

OpenStudy (anonymous):

\[5.00+15.00= 20.00\]

OpenStudy (anonymous):

\[5.00+15.00= 20.00\]

OpenStudy (anonymous):

That is not one of the answers. the answers are x + 0.05x = 15; x = $14.29 x - 0.05x = 15; x = $15.79 0.5x = 15; x = $30 5x = 15; x = $3.00

OpenStudy (anonymous):

then it would be c

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

your welcome